# In-Context Learning State Vector with Inner and Momentum Optimization

### v1

@@ 6.2 Momentum Optimization (RQ2)
As depicted in Table 1, building upon the inner optimized state vector, our proposed momentum optimization algorithm further enhances the effectiveness of the state vector, achieving the best performance on average in all settings. In the zero-shot setting, the momentum optimization boosts the performance of the inner-optimized state vector with an average increase of 1.3% on Llama-2 and 2.4% on GPT-J. In the few-shot setting, state vector with momentum optimization achieves 0.8% average increase on Llama-2 and 1.0% on GPT-J. This reveals the effectiveness of our momentum optimization. With the combination of inner optimization and momentum optimization, our state vector (mom.) surpasses the original variant, showcasing a remarkable improvement of 11.5% for Llama-2 and 8.3% for GPT-J in the zero-shot setting. In the few-shot setting, our state vector (mom.) still outperforms the task vector with a 2.0% improvement for Llama-2 and 2.7% for GPT-J. Furthermore, without inputting demonstration during inference, the state vector (mom.) achieves an impressive 90% ICL performance on Llama-2 and 78% ICL performance on GPT-J. When compared to ICL with the same examples as the demonstration, state vector (mom.) outperforms ICL in both Llama-2 and GPT-J. These improvements verify the effectiveness of our progressive optimization strategy. Note that applying momentum optimization directly to task vectors does not yield average improvements across tasks in our preliminary experiment. We speculate that this inconsistency stems from the poor robustness of the task vectors, which hinders the stable optimization by momentum optimization and leads to poor performance in some tasks.
@@ 6.3 Divide-and-Conquer Aggregation (RQ3)
As illustrated in the Figure 2, both the D&C aggregation and average aggregation exhibit similar trends in both few-shot and zero-shot settings. The performance of both aggregation methods initially falls short of the ICL baseline. However, their performance boosts when examples increases. The initial poor performance can be attributed to the limited number of state vectors. Additionally, although the performance of the D&C aggregation initially falls behind that of the average aggregation, it exhibits a more substantial performance improvement when examples increase, ultimately outperforming average aggregation in the multiple example setting, highlighting the efficiency of D&C aggregation.
@@ 7.3 Qualitative Study
We provide the visualization by Principal Component Analysis (PCA) of the original state vector in the Antonym task and English-French translation task. As depicted in Figure 4, we have three observations: (1) State vectors corresponding to the examples occupying the same position tend to form distinct clusters. This clustering pattern suggests a high degree of similarity among state vectors within each example position, despite different contexts. (2) A notable separation is evident between the state vectors originating from the first example and other position examples. This demarcation implies that ICL may begin to effectively function with a few examples. (3) An interesting trend is observable in the movement of these clusters as the example position increases. This trend may be indicative of an accumulation of task-specific information, where each additional example contributes to a more nuanced understanding of the model. These findings suggest a progressive enhancement in the ability of model to internalize and reflect the subtleties of the task at hand. Moreover, these observations reflect the efficacy of momentum optimization to leverage the observed clustering trend.
@@ Appendix A Implementation Details
We employ Llama-2-7B and GPT-J-6B as our LLMs, chosen for their moderate model sizes, open-source and capability for ICL. Our method is orthogonal to the choice of transformer-based decoder-only autoregressive LLMs. In this paper, we employ a 10-shot as the default ICL setting, and use “” as the separate token similar to previous works. We tried other tokens but no significant difference. All the experiments are reported over 5 random seeds. The inference mechanism with state vector we describe in §4.1 has a key hyper-parameter (i.e.the layer ). Previous studies [Hendel et al., 2023] have shown that the choice of has influence on performance. We find the best layer for different tasks via the accuracy on the development set. For the inner optimization in §4.2, we choose the last seven state vectors to optimize. This is because the early state vectors yield subpar performance, primarily due to limitations in the available examples. For the momentum optimization, we choose 0.5 as the retention rate for historical momentum from the options of 0.25, 0.5 and 0.75. We run all the experiments on NVIDIA A100 80G GPUs.
@@ Appendix B More Details about Baseline
The task vector is designed to extract the ICL function from a specific layer’s hidden state within the transformer model. This is achieved by directly replacing the corresponding hidden state during inference for intervention. On the other hand, Todd et al. [2023] first extract the ICL function from the output activations across all attention heads in all transformer layers. These activations are then prioritized based on their causal effect, quantified by the variance in the model’s output space with or without individual activation interventions. The mathematical average of the top 10 causal effect activations is the function vector, which is subsequently added to the hidden state of a specific layer during the inference stage.
In contrast to these methods, our approach for state vector extraction focuses on procuring the ICL procession state from the output activations of the attention heads within the first layers. During inference, we replace the corresponding activations with optimized ones. While functionally equivalent to the forward process of the task vector when disregarding state vector optimization (i.e., the vanilla state vector), our approach offers enhanced mechanical explainability. This is attributable to its motivation from the dual form of in-context learning and gradient decay, as explicated in previous work [Dai et al., 2023, Natan et al., 2023]. Furthermore, inspired by the dual form, we focus on the further optimization process. On the other hand, unlike the function vector which extracts activations based on the causal effects resulting from individual interventions, our method is rooted in the underlying mechanisms of ICL. This strategy not only improves mechanical explainability but also demonstrates greater performance as evidenced by extensive experiments. And experiments also show notably poor performance of the function vector on certain knowledge-based datasets, such as Person-Occupation.
@@ Appendix C More Details about Datasets
AG News [Zhang et al., 2015] contains 7,600 instances. Each instance contains the news headlines and the first few sentences of an article as input, and output corresponding labels include Business, Science, Sports, and World.
@@ Appendix E Natural Text Completions
In this study, we evaluate the effectiveness of the momentum optimized state vector on natural text completions. Given a natural text template, we instruct the model to greedily generate 5 tokens with or without intervention in the zero-shot setting. We use exact match accuracy as the metric. Table 3 shows the result of natural text completions on Llama-2. The performance boosts observed with the momentum-optimized state vector on the separate tokens, indicating that it can guide the model to generate answers correctly. We include more examples of natural text completions in the Appendix.
@@ Appendix G Full Result
In this section, we provide the main result of the other six tasks except the main result, and the average performance across all tasks. As shown in Table 5, our inner optimization and momentum optimization effectively enhance the state vector.
Appendix H Qualitative Study
@@
Appendix I Robustness Analysis

### v2

@@ 5.4 Momentum Optimization (RQ2)
As depicted in Table 1, building upon the inner optimized state vector, our proposed momentum optimization algorithm further enhances the effectiveness of the state vector, achieving the best performance on average in all settings. In the zero-shot setting, the momentum optimization boosts the performance of the inner-optimized state vector with an average increase of 1.3% on Llama-2 and 2.4% on GPT-J. In the few-shot setting, state vector with momentum optimization achieves a 0.8% average increase on Llama-2 and 1.0% on GPT-J. This reveals the effectiveness of our momentum optimization. With the combination of inner optimization and momentum optimization, our state vector (mom.) surpasses the original variant, showcasing a remarkable improvement of 11.5% for Llama-2 and 8.3% for GPT-J in the zero-shot setting. In the few-shot setting, our state vector (mom.) still outperforms the task vector with a 2.0% improvement for Llama-2 and 2.7% for GPT-J. Furthermore, without inputting demonstration during inference, the state vector (mom.) achieves an impressive 90% ICL performance on Llama-2 and 78% ICL performance on GPT-J. When compared to ICL with the same examples as the demonstration, state vector (mom.) outperforms ICL in both Llama-2 and GPT-J. These improvements verify the effectiveness of our progressive optimization strategy. Note that applying momentum optimization directly to task vectors does not yield average improvements across tasks in our preliminary experiment. We speculate that this inconsistency stems from the poor robustness of the task vectors, which hinders the stable optimization by momentum optimization and leads to poor performance in some tasks.
@@ 5.5 Divide-and-Conquer Aggregation (RQ3)
As illustrated in the Figure 2, both the D&C aggregation and average aggregation exhibit similar trends in both few-shot and zero-shot settings. The performance of both aggregation methods initially falls short of the ICL baseline. However, their performance boosts when examples increase. The initial poor performance can be attributed to the limited number of state vectors. Additionally, although the performance of the D&C aggregation initially falls behind that of the average aggregation, it exhibits a more substantial performance improvement when examples increase, ultimately outperforming average aggregation in the multiple example setting, highlighting the efficiency of D&C aggregation.
@@ 6.3 Qualitative Study
We provide the visualization by Principal Component Analysis (PCA) of the original state vector in the Antonym, English-French and Product-Company task. As depicted in Figure 4, we have three observations: (1) State vectors corresponding to the examples occupying the same position tend to form distinct clusters. This clustering pattern suggests a high degree of similarity among state vectors within each example position, despite different contexts. (2) A notable separation is evident between the state vectors originating from the first example and other position examples. This demarcation implies that ICL may begin to effectively function with a few examples. (3) An interesting trend is observable in the movement of these clusters as the example position increases. This trend may be indicative of an accumulation of task-specific information, where each additional example contributes to a more nuanced understanding of the model. These findings suggest a progressive enhancement in the ability of model to internalize and reflect the subtleties of the task at hand. Moreover, these observations reflect the efficacy of momentum optimization to leverage the observed clustering trend.
@@ Appendix A Implementation Details
In this paper, we use random sampling to create subsets for each dataset. Each subset consists of 10 instances for demonstrations and one instance for a dummy query since we employ a 10-shot as the default ICL setting. The remaining instances are split into test and development sets with a 7:3 ratio. For experiments with multiple examples, we sample 100 instances instead of 10. We use “” as the separate token similar to previous works. We tried other tokens but no significant difference. All the experiments are reported over 5 random seeds. The inference mechanism with state vector we describe in §4.1 has a key hyper-parameter (i.e.the layer ). Previous studies [Hendel et al., 2023] have shown that the choice of has an influence on performance. We find the best layer for different tasks via the accuracy of the development set. For the inner optimization in §4.2, we choose the last seven state vectors to optimize. This is because the early state vectors yield subpar performance, primarily due to limitations in the available examples. For the momentum optimization, we choose 0.5 as the retention rate for historical momentum from the options of 0.25, 0.5 and 0.75. We run all the experiments on a single NVIDIA A100 80G GPUs. Each of our experiments consumes between 10 minutes to 8 hours of GPU time, depending on the dataset.
@@ Appendix B More Details about Baseline
The task vector is designed to extract the ICL function from a specific layer’s hidden state within the transformer model. This is achieved by directly replacing the corresponding hidden state during inference for intervention. On the other hand, Todd et al. [2023] first extracts the ICL function from the output activations across all attention heads in all transformer layers. These activations are then prioritized based on their causal effect, quantified by the variance in the model’s output space with or without individual activation interventions. The mathematical average of the top 10 causal effect activations is the function vector, which is subsequently added to the hidden state of a specific layer during the inference stage.
In contrast to these methods, our approach for state vector extraction focuses on procuring the ICL procession state from the output activations of the attention heads within the first layers. During inference, we replace the corresponding activations with optimized ones. While functionally equivalent to the forward process of the task vector when disregarding state vector optimization (i.e., the vanilla state vector), our approach offers enhanced mechanical explainability. This is attributable to its motivation from the dual form of in-context learning and gradient decay, as explicated in previous work [Dai et al., 2023, Natan et al., 2023]. Furthermore, inspired by the dual form, we focus on the further optimization process. On the other hand, unlike the function vector which extracts activations based on the causal effects resulting from individual interventions, our method is rooted in the underlying mechanisms of ICL. This strategy not only improves mechanical explainability but also demonstrates greater performance as evidenced by extensive experiments. Experiments also show notably poor performance of the function vector on certain knowledge-based datasets, such as Person-Occupation.
@@ Appendix C More Details about Datasets
AG News [Zhang et al., 2015] contains 7600 instances. Each instance contains the news headlines and the first few sentences of an article as input, and output corresponding labels include Business, Science, Sports, and World.
@@ Appendix E Natural Text Completions
In this study, we evaluate the effectiveness of the momentum optimized state vector on natural text completions. Given a natural text template, we instruct the model to greedily generate 5 tokens with or without intervention in the zero-shot setting. We use exact match accuracy as the metric. Table 3 shows the result of natural text completions on Llama-2. The performance boosts observed with the momentum-optimized state vector on the separate tokens indicate that it can guide the model to generate answers correctly. We include more examples of natural text completions in the Appendix.
@@ Appendix G Full Result
In this section, we provide the additional result with llama-2-7B GPT-J model. We first present the main result of optimization on the other six tasks except the main result, and the average performance across all tasks. As shown in Table 5, our inner optimization and momentum optimization effectively enhance the state vector.
Moreover, we provide the result of state vector aggregation on two additional datasets. As shown in Figure 6, the trends of both D$C and average aggregation follow a similar pattern to the main result shown in Figure 2 as the number of examples increases, illustrating the effectiveness of our aggregation methods.
Appendix H Result on Larger Model
| Model | Method | Antonym | English-French | Person-Instrument | Average | |
| Llama-2-13B | Zero-shot | Regular | 1.2 0.7 | 0.2 0.2 | 0.0 0.0 | 0.5 |
| Function vector | 47.1 1.6 | 23.24.3 | 0.1 0.1 | 23.5 | ||
| Task vector | 46.0 2.4 | 43.17.2 | 58.26.3 | 49.1 | ||
| State vector (inn.) | 47.0 1.2 | 50.51.9 | 66.63.1 | 54.7 | ||
| State vector (mom.) | 47.9 1.1 | 55.93.4 | 68.52.0 | 57.4 | ||
| Few-shot | ICL baseline | 67.0 0.1 | 74.51.3 | 75.00.2 | 72.2 | |
| Function vector | 65.7 1.7 | 75.22.6 | 72.20.4 | 71.3 | ||
| Task vector | 64.8 1.2 | 70.53.5 | 70.63.1 | 68.6 | ||
| State vector (inn.) | 65.5 0.8 | 75.81.6 | 77.01.3 | 72.8 | ||
| State vector (mom.) | 65.9 0.7 | 75.60.4 | 78.61.1 | 73.4 | ||
In this section, we provide the optimization and aggregation results on the larger model. Here we choose Llama-2-13B as its memory requirements suit our hardware conditions. We present the result of the optimization method on three representative datasets shown in Table 6, and the result of the aggregation method on four representative datasets shown in Figure 7. The result shows that our inner and momentum optimization and D&C aggregation method could also benefit the state vector on the larger model setting.
Appendix I Qualitative Study
@@
Appendix J Robustness Analysis
@@
Appendix K Limitation
The definition of state vectors is contingent upon specific assumptions and lacks a rigorous theoretical foundation, which may impact its generalizability and reliability across different NLP tasks. Additionally, the experiments were conducted on a limited scale with moderate-sized models and datasets. These constraints may affect the applicability of the results to larger models or more complex datasets. Further research will explore these aspects to establish a more robust validation of the proposed methods.